The table below gives the list price and the number of bids received for five randomly selected items sold through online auctions. Using this data, consider the equation of the regression line, yˆ=b0+b1xy^=b0+b1x, for predicting the number of bids an item will receive based on the list price. Keep in mind, the correlation coefficient may or may not be statistically significant for the data given. Remember, in practice, it would not be appropriate to use the regression line to make a prediction if the correlation coefficient is not statistically significant.

Price in Dollars 2424 2727 3232 3434 3939
Number of Bids 22 55 66 77 88

Table

Copy Data

Step 1 of 6 :  

Find the estimated slope. Round your answer to three decimal places.
Answer #1

Step 1:

So slope is

Sum of X = 156
Sum of Y = 28
Mean X = 31.2
Mean Y = 5.6
Sum of squares (SSX) = 138.8
Sum of products (SP) = 51.4

Regression Equation = ŷ = bX + a

b = SP/SSX = 51.4/138.8 = 0.370

Step 2: a = MY - bMX = 5.6 - (0.37*31.2) = -5.954

Step 3: ŷ = 0.370X - 5.954